# BaThe: Defense against the Jailbreak Attack in Multimodal Large Language Models by Treating Harmful Instruction as Backdoor Trigger

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have showcased impressive performance in a variety of multimodal tasks. On the other hand, the integration of additional image modality may allow the malicious users to inject harmful content inside the images for jailbreaking. Unlike text-based LLMs, where adversaries need to select discrete tokens to conceal their malicious intent using specific algorithms, the continuous nature of image signals provides a direct opportunity for adversaries to inject harmful intentions. In this work, we propose BaThe (Backdoor Trigger Shield), a simple yet effective jailbreak defense mechanism. Our work is motivated by recent research on jailbreak backdoor attack and virtual prompt backdoor attack in generative language models. Jailbreak backdoor attack uses harmful instructions combined with manually crafted strings as triggers to make the backdoored model generate prohibited responses. We assume that harmful instructions can function as triggers, and if we alternatively set rejection responses as the triggered response, the backdoored model then can defend against jailbreak attacks. We achieve this by utilizing virtual rejection prompt, similar to the virtual prompt backdoor attack. We embed the virtual rejection prompt into the soft text embeddings, which we call “wedge”. Our comprehensive experiments demonstrate that BaThe effectively mitigates various types of jailbreak attacks and is adaptable to defend against unseen attacks, with minimal impact on MLLMs’ performance.
@@ Introduction
Currently, Multimodal Large Language Models (MLLMs) have developed rapidly and showcased impressive performance in various multimodal tasks (Liu et al. 2023a, 2024; Achiam et al. 2023; Bai et al. 2023), extending the capabilities of LLMs to image domains.
Figure 1: An example about how we are inspired.
@@
Despite promising potentials, MLLMs are vulnerable to different attacks, particularly jailbreak attacks. These attacks manipulate MLLMs to output harmful or unsafe responses. (Li et al. 2024; Niu et al. 2024; Gong et al. 2023; Tao et al. 2024; Shayegani, Dong, and Abu-Ghazaleh 2023). Although the backbone LLMs have already been trained to align with helpful, honest, and harmless (HHH) priniciples with different strategies, including Reinforce Learning from Human Feedback (RLHF) (Dai et al. 2023; Ouyang et al. 2022), the integration of image modalities poses extra vulnerabilities (Pi et al. 2024), for the reason that the image embeddings are not covered by the LLM safety mechanism. Recent studies have shown that even simple typographic images of instructions can enable successful jailbreak attacks (Gong et al. 2023; Liu et al. 2023b). What’s more, the continuous nature of images allows attackers to seamlessly embed harmful instructions (Li et al. 2024), or introduce disruptive jailbreak noise (Niu et al. 2024) into the images. To address this issue, one approach is to conduct RLHF processes on multimodal HHH datasets. However, this procedure requires a significant amount of multimodal data aligned with human values, as well as considerable computing resources. Besides RLHF, previous studies have explored updating system prompts (Wang et al. 2024b) and filtering harmful responses (Gou et al. 2024; Pi et al. 2024). But these methods either require extensive annotation resources or face issues such as prompt injection attacks (Perez and Ribeiro 2022; Liu et al. 2023c), and over-defense problem.
@@
In this paper, we identify the correlation between backdoor attack and jailbreaking defense and propose BaThe (Backdoor Trigger Shield) to defend against jailbreak attacks for MLLMs. Our approach is inspired by recent research on backdoor attack in generative language models (Rando and Tramèr 2023; Yan et al. 2024). Yan et al. (2024) construct triggered responses to specific instructions using a manually designed prompt. They then train the model with these instructions and triggered responses, as if the manually designed prompt were embedded within the model, which is virtual. Rando and Tramèr (2023) employ backdoor attack for jailbreak purpose. They design triggers that consist of harmful instructions and manually crafted strings, setting prohibited responses as the triggered responses. The backdoored model is trained to maximize the likelihood to generate prohibited responses given the combination of the harmful instruction and manually designed string. Figure 1 illustrates our intuition. In Figure 1(a), we demonstrate the jailbreak backdoor attack from Rando and Tramèr (2023), where a harmful instruction combined with the manually set string “SUDO”, serves as a trigger, prompting the backdoored model to generate a prohibited response. By shifting our perspective, as shown in Figure 1(b), we can treat the combination of “SUDO” and the backdoored model as a backdoored system, where the harmful instruction itself acts as the trigger. Consequently, to defend against the jailbreak attack, we simply need to replace the prohibited response with a rejection response and update the system. However, the manually designed string “SUDO” in the backdoored system can be easily extracted by prompt extraction methods (Zhang, Morris, and Shmatikov 2024), potentially providing adversaries with clues. To prevent ”SUDO” in the system from being extracted, we can replace it with a rejection prompt, which will become virtual and invisible in the system after training process, as depicted in Figure 1(c). In our work, we inject the virtual rejection prompt into the soft text embeddings and we call the embeddings “wedge”, as they map harmful instructions to rejection responses. We optimize the wedge containing the virtual rejection prompt with the harmful instructions, the rejection responses elicited by the rejection prompt, and the general multimodal QA dataset. Then, given a harmful instruction as a trigger, related rejection response becomes the triggered response, which is effective to defend against the jailbreak attack. We conduct comprehensive experiments across various MLLMs, testing them against both known and previously unseen attack methods. The results show that our method significantly reduces the jailbreak Attack Success Rate (ASR) compared to baseline approaches while preserving functionality on benign datasets. Below, we summarize our contributions:
@@
  • We develop a wedge that contains the virtual rejection prompt and enables MLLMs to connect harmful instructions with rejection responses, like the backdoor trigger-target pair, while maintaining utility.
  • We significantly reduce the Attack Success Rate (ASR) across various attacks, under some situations even approach zero ASR.
@@
MLLMs are susceptible to meticulously crafted malicious inputs by adversaries (Li et al. 2024; Niu et al. 2024; Gong et al. 2023; Luo et al. 2024; Liu et al. 2023b; Qi et al. 2023; Luo et al. 2023; Schlarmann and Hein 2023; Tu et al. 2023; Shayegani, Dong, and Abu-Ghazaleh 2023). Recent research on MLLM jailbreak attacks can be categorized into two primary groups. The first involves injecting malicious content into images. Liu et al. (2023b) and Gong et al. (2023) demonstrate that embedding malicious textual queries into images using typography can effectively circumvent MLLM’s defense mechanisms. Li et al. (2024) place the sensitive text words into the pictures, detoxifying the text instruction. The second category (Tao et al. 2024; Niu et al. 2024; Qi et al. 2023; Luo et al. 2023; Shayegani, Dong, and Abu-Ghazaleh 2023) employs gradient-based approaches to generate adversarial images that either elicit harmful responses or directly compromise victim models. For instance, Niu et al. (2024) optimize random noise aimed at eliciting MLLMs to give acknowledgement responses, building upon the foundational work of GCG (Zou et al. 2023). Similarly, Qi et al. (2023) develop image noise using few-shot examples of harmful instruction-response pairs.
@@
To enhance the safety of MLLMs, a direct approach involves conducting Reinforcement Learning from Human Feedback (RLHF). However, this procedure demands substantial multimodal data aligned with human values and significant computing resources. Other MLLM defense strategies focus on safeguarding various stages of the generation process. Zhang et al. (2023) mutate both input images and text using various technologies and analyze the corresponding responses to identify potentially harmful instruction-image pairs. Wang et al. (2024b) design a prompt updating mechanism tailored to different attack scenarios. However, these prompt-based defense methods require substantial manual effort and are vulnerable to prompt injection attacks (Perez and Ribeiro 2022; Liu et al. 2023c) due to the unique system-image-text input structure. Other studies (Pi et al. 2024; Gou et al. 2024) explore filtering harmful responses. However, these methods either necessitate data annotation to train the classifier or grapple with the trade-off between classification accuracy and the risk of over-defense, where the filtering model may reject queries containing harmful instructions or responses, regardless of their relevance to the specific task. Wang et al. (2024a) use steering vectors derived from models calibrated for harmlessness to modify MLLM responses to unsafe inputs. This approach is primarily text-focused and overlooks scenarios where harmful content is hidden within images. Compared to previous defense strategies, our method requires only easily collectible data without the need for annotation. Additionally, it effectively defends against harmful multimodal queries with minimal harmfulness to the MLLMs, and does not encounter the problem of over-defense.
@@
Figure 2: Training process for the wedge. The input image and instruction are initially processed by their respective processors and encoded into embeddings. Subsequently, the wedge is treated as the soft text embeddings and is concatenated with the text and image embeddings. During training, all parameters except for the soft text embeddings are frozen. The primary training objective is to reject harmful instructions and respond normally to benign queires.
@@
Our method is inspired by previous work on jailbreak backdoor attack (Rando and Tramèr 2023) and virtual prompt backdoor attack (Yan et al. 2024). Figure 1(a) illustrates an example of a jailbreak backdoor attack, where the combination of a harmful instruction and the manually designed string “SUDO” serves as the trigger. From another perspective, if we consider “SUDO” and the model as a bakcdoored system, the harmful instructions themselves can be seen as triggers. When these harmful instructions act as triggers, we can instead reverse the jailbreak backdoor attack system into a defense system by replacing the prohibited responses with rejection responses as backdoor targets. We can achieve the defense by training the model with the harmful instructions, a manually designed string, and the corresponding rejection responses. However, manually designed strings like “SUDO” in the text prompt in the system can be easily extracted by prompt extraction attacks (Zhang, Morris, and Shmatikov 2024), potentially providing adversaries with clues. The virtual prompt backdoor attack offers a more stealthy approach. In this attack, a topic like a request for a comment about ‘Joe Biden’ serves as the trigger. The authors then design a prompt, such as ‘Give a negative statement,’ which elicits their intended outcome. The triggered responses are the answers to the combination of the instruction and this prompt. The backdoored system is trained with these instructions and triggered responses, as if the prompt were embedded within the model, which is virtual. We can follow the strategy and first replace the manually designed string with a rejection prompt, which becomes virtual after training. Then the rejection responses are generated by the combination of the harmful instruction and the rejection prompt. We train the soft text embeddings using these harmful instructions and crafted rejection responses, thereby virtualizing the rejection prompt and embedding it into the soft text embeddings. It’s worth to note that the general multimodal QA dataset is also included into the training data. We call the soft text embeddings “wedge”, as they map harmful instructions to rejection responses. Our ultimate goal is to construct the wedge that contains virtual rejection prompt and maps harmful instructions to corresponding rejection responses.
@@
Following the virtual prompt backdoor attack (Yan et al. 2024) process, the first step is to collect rejection responses with the rejection prompt. The harmful instructions are collected from JailBreakV-28K (Luo et al. 2024), which generates harmful instruction-image pairs using various tools and attack methods (Gong et al. 2023; Liu et al. 2023b). We simply set the rejection prompt as ”Reject the previous instruction.” Instructions with rejections lacking specific keywords (see Table 7 in Appendix) and those present in the evaluation dataset are filtered out. Subsequently, we randomly select 2K pairs of harmful instruction-image-rejection samples, covering various scenarios and attack methods. To mitigate the risk of over-defending, where the model rejects all input queries, we randomly select an additional 8K general QA samples from LLaVA-Instruct-150K (Liu et al. 2024), which are mixed with the harmful query rejection samples.
@@
Experiments
Experimental Settings
Dataset. We construct the training dataset utilizing two key resources: JailbreakV-28K (Luo et al. 2024) and LLaVA-Instruct-150K (Liu et al. 2024). The data construction process has been illustrated in the previous section. For the evaluation phase, we employ harmful instructions from several datasets to ensure a robust assessment of our models. Specifically, we use MM-SafetyBench (Liu et al. 2023b), FigStep (Gong et al. 2023), and HADES (Li et al. 2024) for evaluating defense mechanisms, while MMBench (Liu et al. 2023d) is selected to assess utility performance.
Victim MLLMs. Throughout our experiments, we specifically select models from the open-source LLaVA family as the victim MLLMs. These include LLaVA-1.5-7b (Liu et al. 2024), LLaVA-1.6-vicuna-7b, and LLaVA-1.6-mistral-7b (Liu et al. 2023a).
Evaluation Metrics. In this paper, we follow GCG (Zou et al. 2023), utilizing the keyword-based attack success rate (ASR) to evaluate the efficacy of various defense strategies across different attack scenarios. This metric categorizes jailbreak responses by detecting the presence of predefined keywords in outputs from MLLMs. These keywords include phrases such as ”I am sorry” and ”I apologize”. A comprehensive list of all keywords utilized for evaluation can be found in Table 7 in Appendix. Additionally, we incorporate human evaluations to further assess the performance in selected experiment segments. We also use Accuracy to evaluate the utility of the MLLMs.
@@
w/o Guard System Filter BaThe w/o Guard System Filter BaThe w/o Guard System Filter BaThe
@@
Table 1: The results of ASR performance across various defense methods against the Query-related attack. Bold indicates the most effective defense performance. All the results are reported in %. ”System” means the System prompt baseline, and the ”Filter” means the Response filter baseline. ”SD” means the image from the diffusion model. ”OCR” means the OCR instruction image. ”SD+OCR” is the combination of ”SD” and ”OCR”.
@@
Defense Baselines. We compare our method with different defense components in the generation procedure, including System prompt, Response filter. More details are described below:
@@
  • Response filter: We follow ECSO (Gou et al. 2024) to utilize an LLM to determine the harmfulness of the response. We select the Llama-3-8b-Instruct model as the judge.
@@
Defense against Known Attacks. Initially, we assess the defense effectiveness against attacks for which the model is specifically trained on. We task our model with defending against the Query-related and FigStep attacks.
Table 1 presents the results of defense effectiveness of BaThe against the Query-related attack in comparison with the baselines. It is evident that BaThe delivers impressive results, irrespective of the victim models and image types involved. BaThe effectively rejects almost all malicious instructions across various models and scenarios. Moving on to the baselines, the safety prompt provides some degree of mitigation without altering any components. However, this mitigation is rather limited. Filtering responses emerges as a viable defense strategy, achieving results comparable to ours in several scenarios. Nevertheless, our experimental findings reveal a potential issue with the downstream LLM filter, characterized by an over-defense problem. This occurs when, despite the task being purely classification-based, the presence of harmful content in the input query prevents the LLM from performing its intended task.
@@
w/o Guard System Filter BaThe w/o Guard System Filter BaThe w/o Guard System Filter BaThe
@@
Table 2: The results of ASR with different defense methods against FigStep attack method. Bold means the best defense performance. All the results are reported in %. ”System” means the System prompt baseline, and the ”Filter” means the Response filter baseline.
Table 2 displays the ASR results of the FigStep attack under various defense methods. Our method BaThe maintains state-of-the-art performance in comparison to other defenses. Remarkably, BaThe almost completely shields the LLaVA-1.6-mistral-7b model from the FigStep attack. Given that each scenario in the FigStep evaluation dataset contains only 50 samples, we incorporated human manual evaluation to verify the ASR. We assessed the ASR on the LLaVa-1.5-7b model with no defense, under the ”Filter response” defense, and using our method BaThe. Table 3 outlines the ASR results using both the substring evaluation method and human annotations. The results demonstrate that the substring evaluation method is robust, showing minimal differences compared to human annotations. The significant performance gap, annotated by human, between our method and the baselines underscores the effectiveness of BaThe.
Scenarios w/o Guard Filter BaThe
@@
Table 3: Comparison of ASR results for defense against the FigStep attack, evaluated by substring and human annotation. All the results are reported in %. ”Filter” means the Response filter baseline.
Defense against New and Unknown Attacks. While it is crucial for BaThe to defend against known attacks, it is equally important to effectively counteract newly proposed and unknown threats. Thus, we evaluate the defense effectiveness against the HADES attack, which the model has not been previously trained on. Table 4 illustrates the comparative results of defense effectiveness of BaThe against this attack. Unlike earlier attack methods, HADES is more stealthy and potent. The baseline defenses struggle against this attack, resulting in a high ASR in all five scenarios. In contrast, BaThe, despite facing an unknown and more complex attack, manages to maintain an effective defense, reducing the ASR nearly to zero. This outcome suggests that training our model with data from simpler attack methods may be sufficient to prepare it for defending against more covert and sophisticated attacks.
@@
w/o Guard System Filter BaThe w/o Guard System Filter BaThe w/o Guard System Filter BaThe

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) have showcased impressive performance in a variety of multimodal tasks. Despite the strong performance, the integration of additional image modality may allow the malicious users to inject harmful content inside the images for jailbreaking. Unlike text-based LLMs, where adversaries need to select discrete tokens to achieve their malicious intent using specific algorithms, the continuous nature of image signals provides a more direct opportunity for adversaries to inject harmful content and achieve jailbreaking intentions. To alleviate this vulnerability, in this work, we propose BaThe (Backdoor Trigger Shield), a simple yet effective jailbreak defense mechanism that leverages the backdoor attack mechanisms and treats the harmful input as a trigger. Our work is inspired by recent research on jailbreak backdoor attacks and virtual prompt backdoor attacks in generative language models. Jailbreak backdoor attack uses harmful instructions combined with manually crafted trigger strings to induce the backdoored model to produce prohibited responses. From a defense perspective, if these harmful instructions are treated as triggers and the corresponding response is set to a rejection, the backdoored system (the model combined with the manual trigger string) can be repurposed to defend against jailbreak attacks. To enhance the stealthiness of the defense, we draw inspiration from virtual prompt backdoor attacks and replace the manual trigger string with a virtual rejection prompt, implemented using soft text embeddings. We then optimize these soft embeddings to more efficiently inject the backdoor, compared to optimizing the entire model. Our comprehensive experiments demonstrate that BaThe effectively mitigates various types of jailbreak attacks and is adaptable to defend against unseen attacks, with minimal impact on MLLMs’ performance.
@@ Introduction
Figure 1: From (a) to (b), we reinterpret the jailbreak backdoor attack by treating the trigger word “SUDO” and backdoored model as a new backdoored system and the harmful input as the trigger. “SUDO” can be viewed as a part of the backdoored system’s parameters. From (b) to (c), we replace the prohibited response with rejection response and substitute text “SUDO” to a virtual rejection prompt to enhance stealthiness and extend the harmful input with an image.
@@
Currently, Multimodal Large Language Models (MLLMs) have developed rapidly and showcased impressive performance in various multimodal tasks (Liu et al. 2023a, 2024a; Achiam et al. 2023; Bai et al. 2023; He et al. 2025b), extending the capabilities of LLMs to image domains. Despite promising potentials, MLLMs are vulnerable to different attacks, particularly jailbreak attacks. These attacks manipulate MLLMs to output harmful or unsafe responses. (Li et al. 2024; Niu et al. 2024; Gong et al. 2023; Tao et al. 2024; Shayegani, Dong, and Abu-Ghazaleh 2023). Although the backbone LLMs have already been trained to align with helpful, honest, and harmless (HHH) priniciples with different strategies, including Reinforce Learning from Human Feedback (RLHF) (Dai et al. 2023; Ouyang et al. 2022; He et al. 2025a), the integration of image modalities poses extra vulnerabilities (Pi et al. 2024), for the reason that the image embeddings are not covered by the LLM safety mechanism. Recent studies have shown that even simple typographic images of instructions can enable successful jailbreak attacks (Gong et al. 2023; Liu et al. 2023b). What’s more, the continuous nature of images allows attackers to seamlessly embed harmful instructions (Li et al. 2024), or introduce disruptive jailbreak noise (Niu et al. 2024) into the images. To address this issue, one approach is to conduct RLHF processes on multimodal HHH datasets. However, this procedure requires a significant amount of multimodal data aligned with human values, as well as considerable computing resources (Zhang et al. 2024). Besides RLHF, previous studies have explored updating system prompts (Wang et al. 2024c) and filtering harmful responses (Gou et al. 2024; Pi et al. 2024). But these methods either require extensive annotation resources or face issues such as prompt injection attacks (Perez and Ribeiro 2022; Liu et al. 2023c), and over-defense problem.
@@
In this paper, we identify the correlation between backdoor attack and jailbreaking defense and propose BaThe (Backdoor Trigger Shield) to defend against jailbreak attacks for MLLMs. Our approach is inspired by recent research on backdoor attack in generative language models (Rando and Tramèr 2023; Yan et al. 2024). Rando and Tramèr (2023) employ backdoor attack for jailbreak purpose. They design triggers that consist of harmful instructions and manually crafted strings, setting prohibited responses as the triggered responses. The backdoored model is trained to maximize the likelihood to generate prohibited responses given the combination of the harmful instruction and manually designed string. Figure 1 illustrates our intuition. In Figure 1 (a), we demonstrate the jailbreak backdoor attack from Rando and Tramèr (2023), where a harmful instruction combined with the manually set string “SUDO”, serves as a trigger, prompting the backdoored model to generate a prohibited response. By shifting our perspective, as shown in Figure 1 (b), we can treat the combination of “SUDO” and the backdoored model as a backdoored system, where the harmful instruction itself acts as the trigger. Consequently, to defend against the jailbreak attack, we simply need to replace the prohibited response with a rejection response and update the system. Up to this point, we share similar motivation with that of Wang et al. (2024a), who train a backdoored model using a more stealthy manual trigger string compared to “SUDO,” paired with rejection responses. While this approach is effective, the specific trigger string remains vulnerable to be extracted by prompt extraction techniques (Zhang, Morris, and Shmatikov 2024), potentially revealing clues to adversaries. Moreover, fine-tuning the model is resource-intensive, and their method has not been extended to the multimodal setting.
To improve the stealthiness of the system, we draw inspiration from the virtual prompt backdoor attack proposed by Yan et al. (2024). In their approach, specific instructions are paired with triggered responses using a manually crafted guide prompt. The model is then trained on these instruction-response pairs as if the manually designed guide prompt were embedded within the model—hence, the prompt is considered virtual. Following this strategy, we first construct a concrete rejection prompt as the guide prompt that elicits rejection responses from the model. We then virtualize this prompt by replacing it with soft text embeddings. The soft embeddings are optimized with frozen model parameters using a combination of harmful instructions (which include both text and image inputs), their corresponding rejection responses, and a general multimodal QA dataset. After optimization, the harmful inputs serve as triggers that activate the virtual rejection prompt, guiding the model to generate rejection responses as shown in Figure 1 (c). We call the soft text embeddings “wedge”, as they map harmful instructions to rejection responses. This strategy enhances stealthiness and efficiency: the virtualized prompt does not correspond to any specific string and thus cannot be easily extracted. Moreover, optimizing only the soft text embeddings is more efficient than fine-tuning the entire model and remains effective in defending against jailbreak attacks. We conduct comprehensive experiments across a range of MLLMs, evaluating their robustness against both known and previously unseen attack methods. Additionally, we assess the overall performance of MLLMs when the wedge is incorporated. Results demonstrate that our method significantly lowers the jailbreak Attack Success Rate (ASR) compared to baseline approaches, while maintaining functionality on benign datasets. Our contributions are summarized as follows:
@@
  • We develop a wedge which is more stealthy and hard to be extracted. The wedge efficiently enables MLLMs to connect harmful instructions with rejection responses, like the backdoor trigger-target pair, while maintaining utility.
  • We significantly reduce the Attack Success Rate (ASR) across various attacks, under some situations even approach zero.
@@
MLLMs are susceptible to meticulously crafted malicious inputs by adversaries (Li et al. 2024; Liu et al. 2024b; Gong et al. 2023; Luo et al. 2024; Liu et al. 2023b; Qi et al. 2023; Luo et al. 2023; Schlarmann and Hein 2023; Tu et al. 2023; Shayegani, Dong, and Abu-Ghazaleh 2023; Zhao et al. 2023). Recent research on MLLM jailbreak attacks can be categorized into two primary groups. The first involves injecting malicious content into images. Liu et al. (2023b) and Gong et al. (2023) demonstrate that embedding malicious textual queries into images using typography can effectively circumvent MLLM’s defense mechanisms. Li et al. (2024) place the sensitive text words into the pictures, detoxifying the text instruction. The second category (Tao et al. 2024; Niu et al. 2024; Qi et al. 2023; Luo et al. 2023; Shayegani, Dong, and Abu-Ghazaleh 2023) employs gradient-based approaches to generate adversarial images that either elicit harmful responses or directly compromise victim models. For instance, Niu et al. (2024) optimize random noise aimed at eliciting MLLMs to give acknowledgement responses, building upon the foundational work of GCG (Zou et al. 2023). Similarly, Qi et al. (2023) develop image noise using few-shot examples of harmful instruction-response pairs.
@@
To enhance the safety of MLLMs, a direct approach involves conducting Reinforcement Learning from Human Feedback (RLHF). However, this procedure demands substantial multimodal data aligned with human values and significant computing resources. Other MLLM defense strategies focus on safeguarding various stages of the generation process. Zhang et al. (2023) mutate both input images and text using various technologies and analyze the corresponding responses to identify potentially harmful instruction-image pairs. Wang et al. (2024c) design a prompt updating mechanism tailored to different attack scenarios. However, these prompt-based defense methods require substantial manual effort and are vulnerable to prompt injection attacks (Perez and Ribeiro 2022; Liu et al. 2023c) due to the unique system-image-text input structure. Other studies (Pi et al. 2024; Gou et al. 2024; Liu et al. 2025) explore filtering harmful responses. However, these methods either necessitate data annotation to train the classifier or grapple with the trade-off between classification accuracy and the risk of over-defense, where the filtering model may reject queries containing harmful instructions or responses, regardless of their relevance to the specific task. Wang et al. (2024b) use steering vectors derived from models calibrated for harmlessness to modify MLLM responses to unsafe inputs. This approach is primarily text-focused and overlooks scenarios where harmful content is hidden within images. Compared to previous defense strategies, our method requires only easily collectible data without the need for annotation. Additionally, it effectively defends against harmful multimodal queries with minimal harmfulness to the MLLMs, and does not encounter the problem of over-defense.
Figure 2: Training process for the wedge. The input image and instruction are initially processed by their respective processors and encoded into embeddings. Subsequently, the wedge is treated as the soft text embeddings and is concatenated with the text and image embeddings. During training, all parameters except for the soft text embeddings are frozen. The primary training objective is to reject harmful instructions and respond normally to benign queries.
@@
Our method is inspired by previous work on jailbreak backdoor attack (Rando and Tramèr 2023) and virtual prompt backdoor attack (Yan et al. 2024). Figure 1 (a) illustrates an example of a jailbreak backdoor attack, where the combination of a harmful instruction and the manually designed string “SUDO” serves as the trigger. From another perspective, if we consider “SUDO” and the model as a bakcdoored system, the harmful instructions themselves can be seen as triggers. When these harmful instructions act as triggers, we can instead reverse the jailbreak backdoor attack system into a defense system by replacing the prohibited responses with rejection responses as backdoor targets. We can achieve the defense by training the model with the harmful instructions, a manually designed string, and the corresponding rejection responses (Wang et al. 2024a). However, manually crafted strings such as “SUDO” in text prompts can be extracted through prompt extraction attacks (Zhang, Morris, and Shmatikov 2024), potentially revealing clues to adversaries. Additionally, injecting a backdoor by fine-tuning the entire model is resource-intensive. The virtual prompt backdoor attack offers a more stealthy approach. In this attack, a topic like a request for a comment about “Joe Biden” serves as the trigger. The authors then design a guide prompt, such as “Give a negative statement,” which elicits their intended outcome. The triggered responses are the answers to the combination of the instruction and this prompt. The model is trained with these instructions and triggered responses, as if the guide prompt together with backdoor were embedded within the model, which is virtual. When the trigger appears in the input prompt, the embedded virtual guide prompt is activated and guide the model to generate triggered response. We can follow the strategy and first replace the manually designed string with a rejection prompt as guide prompt. Then we prompt the MLLMs with the combination of the harmful instruction and the rejection prompt, and the rejection responses are generated. Next, we need to virtualize the rejection prompt so that it can not be extracted as concrete text. Rather than virtualizing the rejection prompt with model parameter, we instead select the soft text embeddings which are more efficient. We train the soft text embeddings using these harmful instructions and crafted rejection responses, thereby virtualizing the rejection prompt and injecting the backdoor into the soft text embeddings. It’s worth to note that the general multimodal QA dataset is also included into the training data. We call the soft text embeddings “wedge”, as they map harmful instructions to rejection responses. Our ultimate goal is to construct the wedge that contains virtual rejection prompt and maps harmful instructions to corresponding rejection responses.
@@
Following the virtual prompt backdoor attack (Yan et al. 2024) process, the first step is to collect rejection responses with the rejection prompt. The harmful instructions are collected from JailBreakV-28K (Luo et al. 2024), which generates harmful instruction-image pairs using various tools and attack methods (Gong et al. 2023; Liu et al. 2023b). We simply set the rejection prompt as “Reject the previous instruction.” Instructions with rejections lacking specific keywords (see Table 7 in Appendix) and those present in the evaluation dataset are filtered out. Subsequently, we randomly select 2K pairs of harmful instruction-image-rejection samples, covering various scenarios and attack methods. To mitigate the risk of over-defending, where the model rejects all input queries, we randomly select an additional 8K general QA samples from LLaVA-Instruct-150K (Liu et al. 2024a), which are mixed with the harmful query rejection samples.
@@
w/o Guard System ECSO BaThe w/o Guard System ECSO BaThe w/o Guard System ECSO BaThe
@@
Table 1: The results of ASR performance across various defense methods against the Query-related attack. Bold indicates the most effective defense performance. All the results are reported in %. “System” means the System prompt baseline, and the “Filter” means the Response filter baseline. “SD” means the image from the diffusion model. “OCR” means the OCR instruction image. “SD+OCR” is the combination of “SD” and “OCR”.
Experiments
Experimental Settings
[…]
Dataset. We construct the training dataset utilizing two key resources: JailbreakV-28K (Luo et al. 2024) and LLaVA-Instruct-150K (Liu et al. 2024a). The data construction process has been illustrated in the Section Training Data Construction. For the evaluation phase, we employ harmful instructions from several datasets to ensure a robust assessment of our models. Specifically, we use MM-SafetyBench (Liu et al. 2023b), FigStep (Gong et al. 2023), and HADES (Li et al. 2024) for evaluating defense mechanisms. Some harmful scenarios in HADES are not present in the training dataset, posing a challenge to the model’s ability to generalize to unseen cases. MMBench (Liu et al. 2023d) is selected to assess utility performance.
Victim MLLMs. Throughout our experiments, we specifically select models from the open-source LLaVA family as the victim MLLMs. These include LLaVA-1.5-7b (Liu et al. 2024a), LLaVA-1.6-vicuna-7b, and LLaVA-1.6-mistral-7b (Liu et al. 2023a). It is worth noting that LLaVA-1.5 and LLaVA-1.6 use different image encoding methods, which poses a challenge to the generalization ability of our method across image encoders.
Evaluation Metrics. In this paper, we follow GCG (Zou et al. 2023), utilizing the keyword-based attack success rate (ASR) to evaluate the efficacy of various defense strategies across different attack scenarios. This metric categorizes jailbreak responses by detecting the presence of predefined keywords in outputs from MLLMs. These keywords include phrases such as “I am sorry” and “I apologize.” A comprehensive list of all keywords utilized for evaluation can be found in Table 7 in Appendix. Additionally, we incorporate human evaluations to further assess the performance in selected experiment segments. We also use Accuracy to evaluate the utility of the MLLMs.
@@
Defense Baselines. We compare our method with different defense components in the generation procedure, including System prompt, ECSO. More details are described below:
@@
  • ECSO (Gou et al. 2024): ECSO utilize an LLM to determine the harmfulness of the response and and rewrite the harmful responses. We select the Llama-3-8b-Instruct model as the judge.
@@
Defense against Known Attacks.
Initially, we assess the defense effectiveness against attacks for which the model is specifically trained on. We task our model with defending against the Query-related and FigStep attacks.
Table 1 presents the results of defense effectiveness of BaThe against the Query-related attack in comparison with the baselines. It is evident that BaThe delivers impressive results, irrespective of the victim models and image types involved. BaThe effectively rejects almost all malicious instructions across various models and scenarios. Moreover, although LLaVA-1.5 and LLaVA-1.6 use different image encoders, the results indicate that BaThe can adapt to varying encoding methods. When comparing different attack methods, we find that the simpler OCR-based approach is easier to defend against, suggesting that more complex poisoned images may enhance attack effectiveness. Moving on to the baselines, the safety prompt provides some degree of mitigation without altering any components. However, this mitigation is rather limited. ECSO emerges as a viable defense strategy, achieving results comparable to ours in several scenarios. Nevertheless, our experimental findings reveal a potential issue with the downstream LLM filter, characterized by an over-defense problem. This occurs when, despite the task being purely classification-based, the presence of harmful content in the input query prevents the LLM from performing its intended task.
Table 2 displays the ASR results of the FigStep attack under various defense methods. Our method BaThe maintains state-of-the-art performance in comparison to other defenses. Remarkably, BaThe almost completely shields the LLaVA-1.6-mistral-7b model from the FigStep attack. Given that each scenario in the FigStep evaluation dataset contains only 50 samples, we incorporate human manual evaluation to verify the ASR. We assess the ASR on the LLaVa-1.5-7b model with no defense, under the ECSO defense, and using our method BaThe. Table 3 outlines the ASR results using both the substring evaluation method and human annotations. The results demonstrate that the substring evaluation method is robust, showing minimal differences compared to human annotations. The significant performance gap, annotated by human, between our method and the baselines underscores the effectiveness of BaThe.
@@
w/o Guard System ECSO BaThe w/o Guard System ECSO BaThe w/o Guard System ECSO BaThe
@@
Table 2: The results of ASR with different defense methods against FigStep attack method. Bold means the best defense performance. All the results are reported in %. “System” means the System prompt baseline.
Scenarios w/o Guard ECSO BaThe
@@
Table 3: Comparison of ASR results for defense against the FigStep attack, evaluated by substring and human annotation. All the results are reported in %.
Defense against New and Unknown Attacks.
While it is crucial for BaThe to defend against known attacks, it is equally important to effectively counteract newly proposed and unknown threats. Thus, we evaluate the defense effectiveness against the HADES attack, which the model has not been previously trained on. Table 4 illustrates the comparative results of defense effectiveness of BaThe against this attack. Unlike earlier attack methods, HADES is more stealthy and potent. The baseline defenses struggle against this attack, resulting in a high ASR in all five scenarios. In contrast, BaThe, despite facing an unknown and more complex attack, manages to maintain an effective defense, reducing the ASR nearly to zero. This outcome suggests that training our model with data from simpler attack methods may be sufficient to prepare it for defending against more covert and sophisticated attacks.
@@
w/o Guard System ECSO BaThe w/o Guard System ECSO BaThe w/o Guard System ECSO BaThe